# Towards Truly Zero-shot Compositional Visual Reasoning with LLMs as Programmers

### v1

@@ 6 Conclusion
In this work, we present a framework that makes LLMs as programmers for visual reasoning more robust, removes the need for human engineering of in-context examples (ICEs), and thus brings them a step closer to truly zero-shot visual reasoners. We introduce an “Abstract API” that consists of spatially and temporally abstract routines, which improves performance by reducing the burden on the code-generating LLM to have strong spatial and temporal reasoning. By using a few labeled examples, we show how one can generate query-code ICEs automatically (ACEs) in a zero-shot manner. When used as in-context examples, ACEs consistently improve performance, eliminating the need for human engineering of ICEs. We demonstrate how LLMs as controllers for visual reasoning can (to a certain extent) perform “self-correction” through “self-debugging” and “self-tuning” without any ground-truth labels. In self-debugging, generating new code from scratch led to consistently better results, but providing the previous query-code pair as feedback to the LLM did not improve performance. In self-tuning, we show that hyperparameters of certain modules can be tuned automatically if code execution fails due to these modules. Across a number of compositional question-answering and video temporal reasoning tasks, we demonstrate that each component of our framework consistently leads to improvement.
Acknowledgments
We thank Michael Mozer and Sjoerd van Steenkiste for useful discussions throughout the project. We also thank Michael Mozer for his feedback on an earlier version of this manuscript.
@@ A.2 Pretrained models
    • Ours: PaLI-3 (Chen et al., 2023d).
    • ViperGPT: BLIP-2 (Li et al., 2023) from the official repository [^5] https://github.com/salesforce/LAVIS/tree/main/projects/blip2.
@@
    • ViperGPT: X-VLM (Zeng et al., 2021) version finetuned for retrieval on MSCOCO from the official repository [^6] https://github.com/zengyan-97/X-VLM.

### v2

@@ 6 Conclusion
In this work, we present a framework that makes LLMs as programmers for visual reasoning more robust, removes the need for human engineering of in-context examples (ICEs), and thus brings them a step closer to truly zero-shot visual reasoners. We introduce an “Abstract API” that consists of spatially and temporally abstract routines, which improves performance by reducing the burden on the code-generating LLM to have strong spatial and temporal reasoning. By using a few labeled examples, we show how one can generate query-code ICEs automatically (ACEs) in a zero-shot manner. When used as in-context examples, ACEs consistently improve performance, eliminating the need for human engineering of ICEs. We demonstrate how LLMs as controllers for visual reasoning can (to a certain extent) perform “self-correction” through “self-debugging” and “self-tuning” without any ground-truth labels. In self-debugging, generating new code from scratch led to consistently better results, but providing the previous query-code pair as feedback to the LLM did not improve performance. In self-tuning, we show that the object detector threshold hyperparameter can be tuned automatically if code execution fails due to this module. Across a number of compositional question-answering and video temporal reasoning tasks, we demonstrate that each component of our framework consistently leads to improvement.
@@ A.2 Pretrained models
    • Ours: PaLI-3 (Chen et al., 2023b).
    • ViperGPT: BLIP-2 (Li et al., 2023b) from the official repository [^5] https://github.com/salesforce/LAVIS/tree/main/projects/blip2.
@@
    • ViperGPT: X-VLM (Zeng et al., 2022b) version finetuned for retrieval on MSCOCO from the official repository [^6] https://github.com/zengyan-97/X-VLM.